Which of the following statistics are unbiased estimators of population parameters? Choose the correct answer below. Select all that apply. □ A. Sample range used to estimate a population range. □ B. Sample variance used to estimate a population variance □ C. Sample proportion used to estimate a population proportion. D. Sample mean used to estimate a population mean □ E. Sample standard deviation used to estimate a population standard deviation. □ F. Sample median used to estimate a population median
Answer #1

An estimator T is called unbiased estimator for population parameter 0 if Ε(T )-θ,νθ. The correct answers are as follows: B.Now, to show that sample variance is the unbiased estimator of population variance, that is, E(s³)= o². Begin with L.H.S. andNow substitute E(X²) and E(X;)in E(s²) as follows: E(-?+ p8) -n E(s*)= п-1 in i=1 [no? +nự – ơ - nự - 1 no² -o²] n-1 п-1 n-1Now to show that the sample proportion is the unbiased estimator of population proportion. Let sample proportion be denoted b